OpenStudy (anonymous):

Fin the distance (-3,-3) (3,4)

OpenStudy (anonymous):

distance formula:\[ \sqrt{(x-x_{0})^{2} + (y-y_{0})^{2}}\]

OpenStudy (anonymous):

i got the square root of 181

OpenStudy (anonymous):

you have sqrt((-3-3)^2+(-3-4)^2)=sqrt(36+49)=sqrt(85)

OpenStudy (anonymous):

\[x _{0}= -3\] \[x = 3\] \[y _{0}= -3\] \[y = 4\] so \[\sqrt{(3-(-3))^{2} + (4-(-3))^{2}}\] = \[\sqrt{36 + 49}\] = \[\sqrt{85}\]

OpenStudy (anonymous):

cause from the points Point A (-3,-3) \[(x_{0}, y_{0})\] and Point B (3,4) \[(x, y)\]